se sales and profits ofoly Enterprises are growing at a rate of 30% per year. At the end Suppo offour years the growth rate will drop to a steady 4%. At the end of year 5, Oly will issue its first dividend in the amount ofGH 0.20 per share. Ifthe required return on Oly is 16%, what is the value of a share of Olys stock? Assume dividends grow at the same rate as earnings afteryear4. a. GHe0.749 b. GHe0.767 c. GHç0.817 d. GHe0.920 e. GHc0.991
Answer #1

value at the end of year 4 = 0.20/(0.16 - 0.04) = 1.6667

value of the share today = 1.6667/1.16^4 = 0.920